### v1

# Rapid-stable Collective Charging and Reverse Suppressing in Many-Body Quantum Batteries

Abstract
Two major challenges in the process of practical development of quantum batteries (QBs) are: how to achieve rapid-stable charging, i.e., the maximum extractable work quantified via ergotropy, and how to identify the reverse mechanism of rapid-stable charging and propose contrary control strategies in many-body quantum systems. This work proposes a strongly coupled many-body collective charging QB scheme interacting with the surrounding environment, analyzes its physical operating mechanisms as a battery system, explores the mechanisms for achieving fast and stable collective charging under strong environmental coupling, and provides a detailed analysis of various physical methods for suppressing the reverse mechanism. This work highlights the novel properties of many-body QB systems and advances the physical research needed to accelerate the practical application of QBs.
PACS numbers
Keywords
Many-body quantum battery; rapid-stable charging;
I Introduction
The rapid development of quantum technologyL. et al. (2016); Xu et al. (2020) has driven the miniaturization of electronic devices, inspiring many innovations in this field. Quantum batteries (QBs)Alicki and Fannes (2012); Andolina et al. (2018a), as atomic-sized energy storage and conversion devices, are a typical example of such innovations. The concept was first introduced by Robert Alicki and his coauthorsAlicki and Fannes (2012) in 2013. Researchers attempt to reveal whether quantum properties can accelerate the charging processRossini et al. (2020a); Ghosh2021FastCO; Santos (2021), and two charging schemes were proposed: a parallel charging scheme, where each subsystem is independently manipulatedGhosh2021FastCO; Santos (2021), and a collective charging scheme, where global unitary operations are applied to the entire Hilbert space of all subsystemsRossini et al. (2020b); Gyhm et al. (2021); Ghosh et al. (2021). And the ultimate goal of all these work is to leverage quantum effects to enhance battery charging processesCampaioli et al. (2018); Farina et al. (2019); Rossini et al. (2019); Julià-Farré et al. (2020), potentially driving technological advancements across various sectors. Such as, Ref.Tirone et al. (2023) investigates work extraction processes from noisy QBs, focusing on the role of nonlocal resources in enhancing efficiency. The results indicate that while noise generally degrades system performance, properly utilizing nonlocal quantum correlations can mitigate its effects and optimize work extraction. An et al.Song et al. (2023) explored the concept of remote charging for QBs, aiming to enhance their performance by addressing the issue of degradation during the charging process, and the authors propose methods to suppress such degradation, thus improving the overall efficiency and longevity of QBs, with potential applications in quantum technologies. A collision model for the charging of a QB was presented by identical nonequilibrium qubit unitsSeah et al. (2021). It shows that when qubits exhibit quantum coherence, interference effects accelerate the energy distribution spread, leading to faster charging compared to classical incoherent protocols. The study also demonstrates that coherent strategies can achieve higher charging power and characterizes the extractable work using the concept of ergotropy.
In common research schemes for optimizing quantum batteries, the focus is mainly on the charging process based on parallelCakmak (2020); Julia-Farre et al. (2018); Le et al. (2017); Zhang et al. (2018) and collective chargingXu et al. (2020); Andolina et al. (2018b); Kamin et al. (2020); Andolina et al. (2018c) under weak coupling conditions. These studies explore how to improve energy storage during the charging process or how to enhance energy transfer to the power-consuming devices during the discharge process. In contrast, this work will investigate how collective charging of QBs under strong coupling conditions can achieve rapid and stable charging, as well as the reverse control mechanisms for this process, providing theoretical solutions for the device research of QBs.
The rest of the paper is organized as follows. In Sec.II, a many-body QB model is expressed in a form similar to a two-level structure(TLSs). In Sec.III, based on the concept of the ergotropy of a QB, we focus on establishing the charging and discharging dynamics of many-body QBs. In Sec.IV, the charging and discharging dynamics of a three-bit QB will be analyzed in detail here. And we discuss the rapid-stable charging and the reverse mechanism effects in this part. And in the final, we draw some conclusions in Sec.VI.
II Many-Body Collective Charging QBs model
In this work, to simulate a QB in practical applications, we proposed an one-dimensional many-body QB strongly coupled to its surrounding environment. All batteries are in the same excited state $|1\rangle$ when fully charged, and in the ground state $|i\rangle_{(i=2,3,\cdots n)}$ when entirely depleted, as shown in the model in Fig.1(a). Let $\Omega_{i}(t)_{(i=2,3,\cdots n-1)}$ represent the charging function of each battery, and $g_{i}(\omega_{i})_{(i=2,3,\cdots n-1)}$ denotes the strong coupling between each battery and its surrounding environment. The tunneling effect between individual battery units is represented by the same parameter $T_{e}$. Since the aforementioned one-dimensional many-body QB achieves the functionality of simultaneous charging of the output terminal by multiple batteries, Fig.1(a) can be simply schemed into Fig.1(b). The same excited state $|1\rangle$ in Fig.1(b) represents the identical state of each battery, i.e., the fully charged state. In the following discussion, the charging dynamics behavior, the strong coupling between each battery and its environment, and the tunneling effects between battery units are all considered as factors that influence the parameters of the many-body QB. Based on this physical model, we conduct a physical analysis and establish the following theoretical model.
Figure 1: Parallel (top) versus collective (bottom) charging schemes. (a) Multiple QBs represented by TLSs are placed side by side for parallel charging with quantum tunneling effects between each battery, (b) the simplified collective charging scheme. Where $\Omega_{i}(t)$ is the charging function for the external field charging the $i$h QB. $g_{i}(\omega_{i})$ is the coupling function for the interaction between the $i$-th QB and the external field, and the wavy lines between adjacent QBs represent the quantum tunneling effects existing between the two QBs.
Although the many-body QB system is composed of $n$ battery units, we can still describe the system using the following Hamiltonian:
$\hat{H}=\hat{H}_{s}+\hat{H}_{e}+\hat{H}_{int},$ (1)
Without loss of generality, we still use a two-level system (TLS) to describe the many-body QB model. In Eq. (1), $\hat{H}_{s}$ represents the system Hamiltonian of the QB, which includes the charging behavior described by the charging function $\Omega_{j}(t)_{(j=2,3,\cdots n-1)}$ and the tunneling effect between individual battery units in addition to the QB itself, as follows,
$\begin{split}\hat{H}_{s}=&\sum^{n}_{i=1}\epsilon_{i}\hat{\sigma}_{i}+\sum_{j=2}^{n-1}\Omega_{j}(t)|1\rangle\langle{j}|\\ &+\sum_{j=2}^{n-1}T_{e}|j\rangle\langle{j+1}|,\end{split}$ (2)
The level $|1\rangle$ represents the fully charged state of the QB, and $|j\rangle$ denotes the entirely depleted state of each individual battery. All these illustrations can be inferred from Fig. 1. To make our proposed many-body strongly coupled QB model more closely resemble a real battery model, the charging function $\Omega_{j}(t)$ in Eq. (2) should be set in the form of a cosine or sine periodic function. In the Debye model, the environmental Hamiltonian $\hat{H}_{e}$ can be expressed as a collection of quantized harmonic oscillators,
$\hat{H}_{e}=\sum_{k=1}^{{n-1}}\omega_{k}\hat{a}^{\dagger}_{k}\hat{a}_{k}.$ (3)
Where $\omega_{k}$ is the frequency of the $k$-th vibrational mode. $\hat{a}^{\dagger}_{k}$ and $\hat{a}_{k}$ are the creation and annihilation operators of the corresponding mode. The mode frequency $\omega_{k}$ is constrained by the Debye cutoff frequency $\omega_{D}$, typically satisfying $\omega_{k}\leq\omega_{D}$. And $\hat{H}_{int}$ in Eq. (1) represents the interaction Hamiltonian between the QBs and the environment, as follows,
$\hat{H}_{int}=\sum_{k=1}^{{n-1}}g_{k}(\hat{A}_{S}\bigotimes\hat{B}_{k}).$ (4)
Here, $g_{k}$ represents the coupling strength between the system and the environment, which typically depends on the wave vector $\mathbf{k}$. $\hat{A}_{S}$ is an operator in the QBs, representing its coupling with the environment, while the degrees of freedom of the environment are described by $\hat{B}_{k}$. In the Debye model, the environment is associated with phonon-related operators, such as the displacement operator of the $k$-th mode in the environment.
III Charge-discharge dynamics of Many-Body QBs
The performance of a battery is primarily determined by two key factors: first, the ability to store the maximum energy in the shortest time; second, the ability to discharge the energy sufficiently within a specified time, achieving the highest discharge efficiency. To get a good QB, we study its performance, i.e., the energy stored, average charging power, and the extractable work. The energy stored in the QB at time $t$ can be defined as,
$E=Tr[\hat{H}_{s}\hat{\rho}(t)]-Tr[\hat{H}_{s}\hat{\rho}(0)].$ (5)
where $\hat{H}_{s}$ is the Hamiltonian of the QB, $\rho(\tau)$ ($\tau$ = $t$ or 00) is state of the QB at time $\tau$, which is is described by the Redfield master equationCrowder et al. (2023); Becker et al. (2022); Zhao et al. (2016) for strongly coupled systems, as follows,
$\begin{split}\frac{d{\hat{\rho}}(t)}{dt}=&-\frac{i}{\hbar}[\hat{H}_{s}\cdot\hat{\rho}(t)-\hat{\rho}(t)\cdot\hat{H}_{s}]\\ &+\sum_{ij}R_{ij}[2\hat{L}_{ij}\hat{\rho}(t)\hat{L}^{\dagger}_{ij}-\{\hat{L}^{\dagger}_{ij}\hat{L}_{ij},\hat{\rho}(t)\}].\end{split}$ (6)
The first term is the interaction of the QB’s Hamiltonian on its density matrix, which describes the dynamics of the QBs themselves. The environment has a large number of degrees of freedom and is unobservable. We aim to “trace out” the environment by performing a partial trace to obtain the effective density matrix of the system, $\rho(t)$. This requires taking the partial trace of the total density matrix $\rho_{\text{total}}(t)$ over the environment, as follows:
$\rho(t)=\text{Tr}_{E}[\rho_{\text{total}}(t)].$ (7)
The second term in Eq. (6) is the environmental interaction term, which describes the interaction between the QBs and the environment through the Lindblad operator $\hat{L}_{ij}$ and the Redfield tensor $R_{ij}$.
III.1 Redfield tensor
In the Redfield master Eq (6), the Redfield tensor $R_{ij}$ describes the coupling strength between the QBs and the environment. This tensor plays a critical role in the spectral distribution of the environment with its frequency $\omega_{k}$. The Redfield tensor is typically defined as,
$R_{ij}=\sum_{k}J_{ij}(\omega_{k})[1+\coth(\frac{\hbar\omega_{k}}{2k_{B}T})]$ (8)
Figure 2: Redfield tensor vs the environmental frequency $\omega_{k}$ and Debye environmental spectral density $J(\omega_{k})$ with other parameters being $\gamma=2.6\times 10^{-4}Hz,T=300K,\omega_{0}=0.05Hz$.
where $J_{ij}(\omega_{k})$ is the spectral density of the environment, describing the coupling strength between the system and the environment. Commonly used spectral density functions include the Debye modelKubo (1966), the Ohmic modelLeggett et al. (1987), and the Drude-Lorentz modelDrude (1900). It is typically a frequency spectrum distribution, representing the coupling modes between the QBs and the environment. At room temperature, the Redfield tensor varies as a function of the environmental frequency $\omega_{k}$ and Debye environmental spectral density $J(\omega_{k})$=$\frac{\gamma\omega_{k}}{\omega_{0}^{2}+\omega_{k}^{2}}$ was shown in Fig. 2. As shown by Fig. 2, the Debye environmental spectral density $J(\omega_{k})$ takes finite values within the range of environmental frequencies $\omega_{k}$, and it can be noted that both have a weakening effect on the Redfield tensor.
The average charging power at time $t$ is valued by $P_{B}$=$E/t$. From the expressions of QB energy storage and instantaneous power, the synchronization of their variations can be observed. Therefore, we will focus on analyzing the dynamic evolution of the energy stored in the QB.
IV Charge-discharge dynamics of collective three-qubit QBs
To measure the charging and discharging dynamics of many-body QBs, we will take the charging and discharging behavior of a three-qubit QB system (i.e., $n=4$) as an example, focusing on how many-body QBs achieve fast charging and the control strategies for counteracting mechanisms that hinder rapid charging. According to Fig. 1, we can write the Hamiltonian of the tthree-qubit QB system as,
$\hat{H}_{s}=\begin{bmatrix}\varepsilon_{1}&\Omega_{12}(t)&\Omega_{13}(t)&\Omega_{14}(t)\\ \Omega_{21}(t)&\varepsilon_{2}&T_{e}&0\\ \Omega_{31}(t)&T_{e}&\varepsilon_{3}&T_{e}\\ 0&0&T_{e}&\varepsilon_{4}\end{bmatrix}$ (9)
In Eq. (9), the charging function in the Hamiltonian are defined as,
$\Omega_{12}(t)=\Omega_{21}(t)=V\sin(\Omega{t}/\tau)$ (10)
$\Omega_{13}(t)=\Omega_{31}(t)=V[1-\cos(\Omega{t}/\tau)]$ (11)
$\Omega_{14}(t)=V\sin(\Omega{t}/\tau)$ (12)
where $V$ is the amplitude of charging function, $\Omega$ is an integer and $\tau$ is the maximum charging time. $\varepsilon_{2}$ = $\varepsilon_{3}$ = $\varepsilon_{4}$ = $0.25eV$, $\varepsilon_{1}$ = $0.25eV$ + $\Delta E$, and $\Delta E$ represents the energy gap. $T_{e}$ is the tunneling effect between different individual battery units.
V Results and discussions
Next, we will focus on examining the dynamical behavior of energy storage in three parallel QBs. Before starting the numerical calculations, some typical parameters need to be predetermined. They are listed in detail in Tab. 1.
Table 1: Parameters for the parallel three-qubit QBs.
$\Omega$ $\Delta{E}(eV)$ $\omega(Hz)$ $\gamma\times 10^{-7}(s)$ $\omega_{0}(Hz)$ $Te$ $T(k)$ $V(\mu V)$
Fig.3(a) $\setminus$ 1.5 0.085 $2.6$ 0.10 0 300 1.5
Fig.3(b) $1.0\pi$setminus 0.085 $2.6$ 0.12 0 300 1.5
Fig.3(c) $1.0\pi$ 2.75 0.085 $2.6$ 0.12 0 300 $\setminus$
Fig.3(d) $1.0\pi$ 2.75 0.085 $2.6$ 0.12 $\setminus$ 300 1.5
Fig.4(a) $1.0\pi$ 1.5 0.085 $2.6$setminus 0 300 1.5
Fig.4(b) $1.0\pi$ 1.5 $\setminus$ $9.0$ 0.12 0 300 1.5
Fig.4(c) $1.0\pi$ 1.5 0.085 $\setminus$ 0.03 0 300 1.5
Fig.4(d) $1.0\pi$ 1.5 0.143 $9.0$ 0.08 0 $\setminus$ 1.5
[…]
The macroscopic classical charging behavior should reflect a proportional relationship with the amplitude of the charging function. In other words, the larger the amplitude of the charging function, the faster the battery should achieve stable and saturated charging. However, in many-body QB systems in Fig. 3(c), different quantum characteristics emerge: when the amplitude is relatively small, the time to reach stable charging is nearly the same. The egotropy, on the other hand, exhibits a linearly increasing trend. But, when the amplitude is larger, the stable charging behavior is disrupted, rapidly dropping to zero. As can be seen from $V$= $3.5\mu V$, $4.0\mu V$. This is a novel feature observed in the many-body QB model.
Refer to captionRefer to caption
Figure 3: The dynamics of optimal energy storage and ergotropy in the parallel three-qubit QBs. The tuning parameters are (a) $\Omega$, (b) $\Delta{E}$, (c) $V$, and (d) $Te$, respectively.
The tunneling effect between many-body QBs is an important parameter that cannot be ignoredZhong et al. (2021). We observe from Fig. 3(d) that the energy storage exhibits an oscillatory evolution with time. When the tunneling effect , $Te$ is small, the ergotropy rapidly drops to zero. Only when the tunneling effect reaches a certain magnitude can the entire system achieve stable ergotropy, such as $Te$ =0.06, and 0.07. This conclusion suggests that in the investigation to the fabrication of many-body QBs, how to utilize the tunneling effect between individual QBs and achieve ergotropy is a key topic for future research.
Refer to captionRefer to caption
Figure 4: Optimal energy storage of the many-body QB system via parameters (a)$\omega_{0}$, (b)$\omega$, (c)$\gamma$, and (d)$T$.
However, the environmental factors in the control parameters were not mentioned in the above discussion. Next, we will focus on the impact of the environmental spectral density and the coupling between the many-body QB system and the environment on the energy storage performance of the battery system. The characteristic frequency $\omega_{0}$ of the Debye spectrum function is reflected as a control parameter for the energy storage in panel (a) of Fig. 4. As can be seen from Fig. 4(a), it can be seen that a larger $\omega_{0}$ can achieve stable energy output in the process of gradual increase of $\omega_{0}$. The results in Fig. 2 tell us about the negative regulatory of the environmental spectral frequency on the Redfield tensor. While the results in Fig. 4(b) indicate that the environmental spectral frequency is beneficial for achieving stable stored energy. Therefore, it can be seen that the Redfield tensor has a negative impact on the stability of stored energy.
From the expression of the Debye environment spectral density $J(\omega_{k})$ , it can be seen that gamma forms a linear relationship with it. However, in Fig. 4(c), $\gamma$ exhibits a negative effect, indicating that $J(\omega_{k})$ has a negative impact on the stable energy output of the many-body QB. Therefore, the reference significance of $J(\omega_{k})$ in the design of many-body QBs. The result in Fig. 4(d) visually shows the destructive effect of environmental temperatures above 100K on QBs’ performance. When the temperature is between 20K and 100K, the environmental temperature has almost no impact on the dynamic behavior of the QBs.
V.1 Possible experimental realization
QBs are energy-efficient storage devices that operate based on quantum tunneling effectsMondal and Bhattacharjee (2021), quantum entanglementImai et al. (2023), and superconducting materialsDou and Yang (2023). Charging refers to the process of transitioning quantum bits from a low-energy state to a high-energy state, while discharging is the reverse process. They can take various physical forms, such as ions, neutral atoms, or photons. In this work, we have proposed a many-body QB model with a tunable energy gap. Theoretically, any quantum devices with the above-mentioned energy gap could serve as an experimental candidate for this model. However, there is a prerequisite that should be met: the gap and tunneling effect parameters of the individual QBs mentioned are the same. Therefore, the physical materials of the individual QBs forming the many-body QB need to be identical in order to achieve the theoretical results predicted by this model. Both two-level systems with appropriate band gaps and quantum spin systems can serve as potential candidates for the experimental realization of this work.
VI Conclusion
In conclusion, we have proposed a strongly coupled many-body QB scheme that enables rapid-stable charging, and we also explored the suppression of the reverse mechanism. It is found that, conclusions contrary to those of macroscopic batteries have been drawn by analyzing its structure of the many-body QB and the external environment, which provides positive guidance for further experimental research. At the same time, some obtained reverse regulation parameters also provide valuable references for experimental design. The significance of this work in the field of QB research lies in the proposal of a strongly coupled many-body QB model. The theoretical significance is in presenting a digital control model for achieving rapid and stable charging.
Author contributions
S. C. Zhao conceived the idea. Y. F. Yang performed the numerical computations and wrote the draft, and S. C. Zhao did the analysis and revised the paper. Z. R. Zhao participated in part of the discussion.
Data Availability Statement
This manuscript has associated data in a data repository. [Authors’ comment: All data included in this manuscript are available upon resonable request by contaicting with the corresponding author.]
Conflict of Interest
The authors declare that they have no conflict of interest. This article does not contain any studies with human participants or animals performed by any of the authors. Informed consent was obtained from all individual participants included in the study.
Acknowledgements.
We gratefully acknowledge support of the National Natural Science Foundation of China ( Grant Nos. 62065009 and 61565008).

### v2

@@
 Researchers attempt to reveal whether quantum properties can accelerate the charging processRossini et al.
-(2020a); Ghosh2021FastCO; Santos (2021), and two charging schemes were proposed: a parallel charging scheme, where each subsystem is independently manipulatedGhosh2021FastCO; Santos (2021), and a collective charging scheme, where global unitary operations are applied to the entire Hilbert space of all subsystemsRossini et al.
+(2020a); Santos (2021), and two charging schemes were proposed: a parallel charging scheme, where each subsystem is independently manipulatedSantos (2021), and a collective charging scheme, where global unitary operations are applied to the entire Hilbert space of all subsystemsRossini et al.
 (2020b); Gyhm et al.
-(2021); Ghosh et al.
 (2021).
@@
 1, we can write the Hamiltonian of the tthree-qubit QB system as,
-$\hat{H}_{s}=\begin{bmatrix}\varepsilon_{1}&\Omega_{12}(t)&\Omega_{13}(t)&\Omega_{14}(t)\\ \Omega_{21}(t)&\varepsilon_{2}&T_{e}&0\\ \Omega_{31}(t)&T_{e}&\varepsilon_{3}&T_{e}\\ 0&0&T_{e}&\varepsilon_{4}\end{bmatrix}$ (9)
+$\hat{H}_{s}=\begin{bmatrix}\varepsilon_{1}&\Omega_{12}(t)&\Omega_{13}(t)&\Omega_{14}(t)\\ \Omega_{21}(t)&\varepsilon_{2}&T_{e}&0\\ \Omega_{31}(t)&T_{e}&\varepsilon_{3}&T_{e}\\ \Omega_{41}(t)&0&T_{e}&\varepsilon_{4}\end{bmatrix}$ (9)
 In Eq. (9), the charging function in the Hamiltonian are defined as,
@@
 $\Omega_{13}(t)=\Omega_{31}(t)=V[1-\cos(\Omega{t}/\tau)]$ (11)
-$\Omega_{14}(t)=V\sin(\Omega{t}/\tau)$ (12)
+$\Omega_{14}(t)=\Omega_{41}(t)=V\sin(\Omega{t}/\tau)$ (12)
 where $V$ is the amplitude of charging function, $\Omega$ is an integer and $\tau$ is the maximum charging time.
@@
 V.1 Possible experimental realization
-QBs are energy-efficient storage devices that operate based on quantum tunneling effectsMondal and Bhattacharjee (2021), quantum entanglementImai et al.
+QBs are energy-efficient storage devices that operate based on quantum tunneling effectsMondal and Bhattacharjee (2021), quantum entanglementShi et al.
 (2022), and superconducting materialsDou and Yang (2023).